OpenStudy (anonymous):

Evaluate lim_{x rightarrow 0} tan5x div x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \tan5x \div x\]

OpenStudy (anonymous):

Could you show how you got that please?

myininaya (myininaya):

\[\lim_{u \rightarrow 0}\frac{\sin(u)}{u}=1\] Do you know this?

myininaya (myininaya):

This is true by squeeze thm We will need this

myininaya (myininaya):

Write in terms of sine and cosine functions

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(5x)}{\cos(5x)} \cdot \frac{1}{x}\] \[\lim_{x \rightarrow 0}\frac{\sin(5x)}{x} \cdot \frac{1}{\cos(5x)}\]

myininaya (myininaya):